# A joint optimization approach to identifying sparse dynamics using least squares kernel collocation

### v1

@@ 4.2.2 Hidden coordinates
We now consider the problem of identifying dynamics with a hidden variable that is never observed. In this case, the observations alternate between the first two coordinates on each time step, and the third is never observed. A similar problem have been considered in [hamzi-partial] and [Ribera2022-nz]. [hamzi-partial] approach this using the lens of computational graph completion [owhadi-CGC] and kernel learning, though the data was without noise, and the model was considered in discrete time. [Ribera2022-nz].
Figure 6: Results for Lorenz 63 system with an unobserved coordinate. Observations alternate between first two coordinates at sampling rate $\Delta t=0.025$ up to $t=10$, with no knowledge of the third, except that the system is known to be three dimensional. Our algorithm constructs a third, latent coordinate and corresponding dynamics which, when evolved together with the first two coordinates, reproduces the the observed dynamics. Further, we we can effectively extrapolate the time series forward by using our learned dynamics and imputing the value of our artificial coordinate as a new initial condition, generating good predictions for the first two coordinates.
@@
Nevertheless, we can learn an extra latent coordinate; when combined with the first two observed coordinates, these together satisfy a modified ODE that reproduces the dynamics on the observed coordinates. This is demonstrated by taking our state estimate at $t=10$, the last observed data-point, and simulating the learned ODE forward in time to predict the evolution of the first two states. Doing so produces accurate trajectories for four additional cycles, including a lobe switch, before diverging from the true trajectory due to chaos. Note that modeling only the first two states with an ODE is insufficient–the flow would be self-intersecting and cannot satisfy any autonomous ODE without a third coordinate (no chaos in two dimensions). The results can be seen in Figure 6 where the following syste is learned in eq. 46. Because the true hidden state is unidentifiable, we compute the state estimation error restricted to the first two coordinates and obtain $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=6.840\cdot 10^{-3}$. Similarly, since we cannot expect to recover the third coordinate, we omit the coefficient error.
@@
Traditional approaches to handling partially observed systems and modeling time series often enrich the dynamics by inserting time delay coordinates [George2021Decomposing, brunton_chaos_2017, rand_detecting_1981, Bakarji2023-sq, Broomhead1989-bh, Lee2019-wf, Pan2020-nd]. However, that comes with its own challenges, such as selecting the number and spacing of delay coordinates, difficulties in continuous time, especially when observations are not evenly spaced, and introducing bias due to possibly noisy inputs. Our approach has its own drawback of requiring a separate state estimation problem to make predictions, and producing a somewhat uninterpretable third coordinate. We remark that it is possible to include time delay embeddings in our approach as well, essentially turning our collocation method for differential equations into a collocation approach for a delay differential equations–we leave this to future work.
@@ 4.3.1 Van der Pol Oscillator
Figure 8: Left: True Von der Pol oscillator trajectories plotted from multiple initial conditions on phrase portrait using eq. 47. Right: Using learned model eq. 48 to simulate same trajectories provided same initial conditions.
@@ 4.3.2 Coupled Duffing Oscillators
As an example of a larger second order ODE, we consider a ring of $d$ Duffing oscillators with nearest-neighbor coupling. For each coordinate $i=1,\dots,d$, $x_{i}(t)$ satisfies the differential equation
@@
We remark that the nonlinearity has a significant qualitative effect in the dynamics. If we consider a linear version where we drop the cubic restoring force, the dynamics would be governed by the matrix $M:=\mathrm{Diag}((\alpha_{i})_{i=1}^{n})+\mu K$, where $K$ is the graph Laplacian on a cycle. Taking $\mu<0$ causes the coupling term to encourage exponential blow up, and with $\mu$ negative enough relative to $\alpha_{i}$, $M$ will have both positive and negative eigenvalues. This implies unbounded level sets in the Hamiltonian, and divergent trajectories in the linear version. However, the positive cubic restoring terms ($\beta_{i}>0$) are quartic in the Hamiltonian, which guarantees compact level sets, bounded trajectories [teschl2012ordinary].
We considered an example in $d=5$ dimensions with parameters $\mu=-2,\;\alpha_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}[2,4],\;\beta_{i}=2$. We initialized the system with all coordinates at the same position and velocity, $x_{i}(0)=0$ and $\frac{d}{dt}x_{i}(0)=0$ for $i=1,\dots,d$. We took equally spaced measurements of the position values $q_{i}(t)$ with a sampling period $\Delta t=0.2$ between $t=0$ and $t=30$ as the input data, and include additive Gaussian noise of variance $\sigma^{2}=0.04$ which corresponds to a signal to noise ratio of $39.2$.
Trajectory regularization, data weight, collocation penalties and points remain as in the previous example $\lambda=10^{-3},\alpha=1,\beta=10^{5},m=1000$. We took the sparsifier to be STLSQ with threshold 0.1, and ridge regularization parameter 0.01. The feature library includes monomials up to cubic order of $q_{i}$ and $\dot{q_{i}}$, but excludes interaction terms,
$\Phi=\left\{1,x_{1},\frac{d}{dt}x_{1},...,x_{5},\frac{d}{dt}x_{5},q_{0}^{2},\dot{q}_{0}^{2},\dots,q_{4}^{3},\dot{q_{4}}^{3}\right\}.$ (52)
@@
Figure 9: Position observations sampled at rate $\Delta t=0.2$ up to $t=30$ with added noise set to $\sigma^{2}=0.04$ equating to a relative noise of $39.2\%$. Learned and true dynamics from eq. 53 are simulated from $t=30$ to $t=50$.
@@ 4.4 A misspecified model
We considered case where the dynamics are misspecified by fitting polynomials to nonlinear pendulum dynamics. Our least squares relaxation of the dynamics allows some of the unmodeled dynamics to be accommodated as an error term, as in eq. 61. Robustness to misspecification is very useful; in many experimental or designed processes we may have little to no prior knowledge of the true functional form of the underlying dynamics. In these cases, we hope to maintain highly accurate state estimation (despite the misspecified dynamics), while producing a reasonable model for simulating the future.
@@
and fit two models, one that is linear and another that is cubic. Simulated trajectories were created with initial displacement of $x(0)=3$ and no initial velocity $\frac{d}{dt}x(0)=0$. Observations were made with a sampling period of $\Delta t=1$ over the interval $t\in[0,100]$, with Gaussian noise ($\sigma^{2}=0.1$).
@@
Both of these perform very well in state estimation, despite only forming a somewhat crude approximation to the dynamics as evaluated through forward simulation (see Figure 11). The filtering error of the the linear model was $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=1.250\cdot 10^{-1}$ and for the the cubic model, was $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=2.221\cdot 10^{-2}$. The linear model matched the period of the true nonlinear dynamics eq. 54, but did not have enough degrees of freedom to match the amplitude and misrepresented the shapes of the peaks. The cubic model performed quite well, especially in state estimation and matching the shapes of the oscillations, but has a slightly incorrect period.
Figure 11: Results for misspecified linear and cubic models for nonlinear pendulum. Observations sampled at rate $\Delta t=1$ up to $t=100$ with added nosie set to $\sigma^{2}=0.01$. Learned dynamics for both linear and cubic model, eq. 55 and eq. 56, are simulated from $t=100$ to $t=200$ where the linear model successfully captures periodicity of the true trajectory, but fails to capture the correct amplitude, whereas this is flipped for the cubic model.
@@ 4.5 Benchmarking
We now consider two systematic experiments testing the robustness of our method to noise, and comparing primarily to ODR-BINDy[Fung2025-odr].
@@ 4.5.1 Damped Nonlinear Oscillator
We considered increasing noise levels with fixed time length and observation frequency for a nonlinear damped oscillator that was studied in [Fung2025-odr]. The dynamics are given by
@@
We took the initial conditions $\boldsymbol{x}(0)=(2,0)$, and took samples of both variables with sampling period $\Delta t=0.05$ for $t\in[0,10]$. We first considered an example with noise level $\sigma=0.32$. We set trajectory regularization $\lambda=10^{-5}$, data weight $\alpha=10$, collocation penalty $\beta=10^{5}$, and used STLSQ with threshold to $0.08$ and ridge parameter to $100$. This is a surprisingly challenging example, especially with this level of noise–our model does not identify the correct support, but still performs quite well in estimating the state. We obtained $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=6.60\times 10^{-2}$, $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}=2.47\times 10^{-1}$, and the equations
@@
Next, we tested on a sequence of noise levels $\sigma=0.02,0.04,0.08,0.16,0.32,0.64$ and averaged the values of $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}$ and $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}$ over $32$ random seeds. In this case, we found that including ensembling [fasel_ensemble-sindy_2022] in the inner sparsifier greatly improved the consistency of the results for higher noise levels. We compared five models, weak SINDy[Messenger2021a], weak SINDy with ensembling, JSINDy, JSINDy with ensembling, and ODR-BINDy [fasel_ensemble-sindy_2022, Fung2025-odr]. We used PySINDy for ensembling and weak SINDy [Kaptanoglu2022], and the GitHub repository [Fung_ODR-BINDy_2025] associated to [Fung2025-odr] for ODR-BINDy. We also tested with other two-step variants of SINDy, but they all performed worse than weak SINDy, so we omit them. As ODR-BINDy also produces state estimates, we compare the state estimation error between JSINDy, JSINDy with ensembling and ODR-BINDy.
@@
In terms of the coefficient error, and without ensembling the sparsifier, our model performs noticeably worse than ODR-BINDy; maximizing the Bayesian evidence seems to a good model selection criterion. Upon introducing ensembling, our model selection stabilizes and the coefficent errors become comparable. In terms of state estimation, JSINDy both with and without ensembling performs well across varying levels of noise, regardless of the coefficient accuracy, owing to the stabilizing effect of the RKHS regularization and least squares relaxation.
@@
Figure 13: Results for the damped nonlinear oscillator across varying levels of noise. Left: coefficient error $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}$, Right: state estimation error $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}$. In this case, ensembling significantly improves the coefficient error for both weak SINDy and JSINDy. For JSINDy, including ensembling inside of the sparsifier brings the coefficient error from noticably worse than ODR-BINDy to slightly better. Overall, JSINDy performs very well in state estimation across varying levels of noise.
@@ 4.5.2 Lorenz 63 example
We also tested the performance of our method on varying levels of noise and durations of observations with fixed observation frequency on the Lorenz 63 dynamics (40) with results in Figure 14. We applied JSINDy with data weight $\alpha=1$, collocation weight $\beta=10^{5}$, regularization $\lambda=10^{-5}$, and used SR3 as a sparsifier with an L1 penalty with weight 5, and relaxation coefficient $\nu=0.05$ [SR3]. We compare to ensemble weak SINDy and ODR-BINDy. The setup is identical to that of [Fung2025-odr, Sec. 3.2], though we did not exclude the constant function from the library for weak SINDy. ODR-BINDy performed best in variable selection, likely owing to their approach of Bayesian evidence maximization. While also better than JSINDy in coefficient error, the difference is somewhat less pronounced. A much larger difference can be seen between JSINDy and ensemble weak SINDy, as well as the other models shown in [Fung2025-odr, Sec. 3.2].
Figure 14: Experiment across different trajectory lengths and noise levels, averaged over 128 independent trials and $\Delta t=0.01$. (3). Top: Proportion of trials where the exact support was exactly recovered Bottom: Relative coefficient error $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}$ of identified system.

### v2

@@ 4.2.2 Hidden coordinates
We now consider the problem of identifying dynamics with a hidden variable that is never observed. In this case, the observations alternate between the first two coordinates on each time step, and the third is never observed. [76] considered a very similar setup, introduced a variational annealing approach to identify dynamics in the presence of hidden variables, and also tested on variants of the Lorenz dynamics. [35] approached the hidden variable problem using the lens of computational graph completion [69] and kernel learning, though only considered problems in discrete time without noise.
Figure 6: Results for Lorenz 63 system with an unobserved coordinate. Observations alternate between first two coordinates at sampling rate $\Delta t=0.025$ up to $t=10$, with no knowledge of the third, except that the system is known to be three dimensional. Our algorithm constructs a third, latent coordinate and corresponding dynamics which, when evolved together with the first two coordinates, reproduces the observed dynamics. Further, we can effectively extrapolate the time series forward by using our learned dynamics and imputing the value of our artificial coordinate as a new initial condition, generating good predictions for the first two coordinates.
@@
Nevertheless, we can learn an extra latent coordinate; when combined with the first two observed coordinates, these together satisfy a modified ODE that reproduces the dynamics on the observed coordinates. This is demonstrated by taking our state estimate at $t=10$, the last observed data-point, and simulating the learned ODE forward in time to predict the evolution of the first two states. Doing so produces accurate trajectories for four additional cycles, including a lobe switch, before diverging from the true trajectory due to chaos. Note that modeling only the first two states with an ODE is insufficient–the flow would be self-intersecting and cannot satisfy any autonomous ODE without a third coordinate (no chaos in two dimensions). The results can be seen in Figure 6 and eq. 46. Because the true hidden state is unidentifiable, we compute the state estimation error restricted to the first two coordinates and obtain $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=6.840\cdot 10^{-3}$. Similarly, since we cannot expect to recover the third coordinate, we omit the coefficient error.
@@
Traditional approaches to handling partially observed systems and modeling time series often enrich the dynamics by inserting time delay coordinates [61, 30, 13, 89, 6, 11, 52, 70]. However, that comes with its own challenges, such as selecting the number and spacing of delay coordinates, difficulties in continuous time, especially when observations are not evenly spaced, and introducing bias due to possibly noisy inputs. Our approach has its own drawback of requiring a separate state estimation problem to make predictions, and producing a somewhat uninterpretable third coordinate. We remark that it is possible to include time delay embeddings in our approach as well, essentially turning our collocation method for differential equations into a collocation approach for a delay differential equations–we leave this to future work.
@@ 4.3.1 Van der Pol oscillator
Figure 8: Left: True Van der Pol oscillator trajectories plotted from multiple initial conditions on phase portrait using eq. 47. Right: Using learned model eq. 48 to simulate same trajectories provided same initial conditions.
@@ 4.3.2 Coupled Duffing oscillators
As an example of a larger second order ODE, we consider a ring of $d$ Duffing oscillators with nearest-neighbor coupling. For each dimension $i=1,\dots,d$, $x_{i}(t)$ satisfies the differential equation
@@
We remark that the nonlinearity has a significant qualitative effect in the dynamics. If we consider a linear version where we drop the cubic restoring force, the dynamics would be governed by the matrix $M:=\mathrm{Diag}((\alpha_{i})_{i=1}^{n})+\mu K$, where $K$ is the graph Laplacian on a cycle. Taking $\mu<0$ causes the coupling term to encourage exponential blow up, and with $\mu$ negative enough relative to $\alpha_{i}$, $M$ will have both positive and negative eigenvalues. This implies unbounded level sets in the Hamiltonian, and divergent trajectories in the linear version. However, the positive cubic restoring terms ($\beta_{i}>0$) are quartic in the Hamiltonian, which guarantees compact level sets, bounded trajectories [90].
We considered an example in $d=5$ dimensions with parameters $\mu=-2,\;\alpha_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}[2,4],\;\beta_{i}=2$. We initialized the system with all coordinates at the same position and velocity, $x_{i}(0)=1$ and $\frac{d}{dt}x_{i}(0)=0$ for $i=1,\dots,d$. We took equally spaced measurements of the position values $x_{i}(t)$ with a sampling period $\Delta t=0.2$ between $t=0$ and $t=30$ as the input data, and include additive Gaussian noise of variance $\sigma^{2}=0.04$ which corresponds to a signal to noise ratio of $39.2$.
Trajectory regularization, data weight, collocation penalties and points remain as in the previous example $\lambda=10^{-3},\alpha=1,\beta=10^{5},m=1000$. We took the sparsifier to be STLSQ with threshold 0.1, and ridge regularization parameter 0.01. The feature library includes monomials up to cubic order of $x_{i}$ and $\frac{d}{dt}x_{i}$, but excludes interaction terms,
$\Phi=\left\{1,x_{1},\frac{d}{dt}x_{1},...,x_{5},\frac{d}{dt}x_{5},x_{1}^{2},(\frac{d}{dt}x_{1})^{2},\dots,x_{5}^{3},(\frac{d}{dt}x_{5})^{3}\right\}.$ (52)
@@
Figure 9: Position observations sampled at rate $\Delta t=0.2$ up to $t=30$ with added noise set to $\sigma^{2}=0.04$ equating to a relative noise of $2.55\%$. Learned and true dynamics from eq. 53 are simulated from $t=30$ to $t=50$.
@@ 4.4 A misspecified model
We considered a case where the dynamics are misspecified by fitting polynomials to nonlinear pendulum dynamics. Our least squares relaxation of the dynamics allows some of the unmodeled dynamics to be accommodated as an error term, as in eq. 61. Robustness to misspecification is very useful; in many experimental or designed processes we may have little to no prior knowledge of the true functional form of the underlying dynamics. In these cases, we hope to maintain highly accurate state estimation (despite the misspecified dynamics), while producing a reasonable model for simulating the future.
@@
and fit two models, one that is linear and another that is cubic. Simulated trajectories were created with initial displacement of $x(0)=3$ and no initial velocity $\frac{d}{dt}x(0)=0$. Observations were made with a sampling period of $\Delta t=1$ over the interval $t\in[0,100]$, with Gaussian noise ($\sigma^{2}=0.01$).
@@
Both of these perform very well in state estimation, despite only forming a somewhat crude approximation to the dynamics as evaluated through forward simulation (see Figure 11). The filtering error of the linear model was $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=1.250\cdot 10^{-1}$ and for the cubic model, was $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=2.221\cdot 10^{-2}$. The linear model matched the period of the true nonlinear dynamics eq. 54, but did not have enough degrees of freedom to match the amplitude and misrepresented the shapes of the peaks. The cubic model performed quite well, especially in state estimation and matching the shapes of the oscillations, but has a slightly incorrect period.
Figure 11: Results for misspecified linear and cubic models for nonlinear pendulum. Observations were sampled at a rate $\Delta t=1$ up to $t=100$ with added noise with variance $\sigma^{2}=0.01$. Learned dynamics for both linear and cubic model, eq. 55 and eq. 56, are simulated from $t=100$ to $t=200$ where the linear model successfully captures periodicity of the true trajectory, but fails to capture the correct amplitude, whereas this is flipped for the cubic model.
@@ 4.5 Benchmarking
We now consider two systematic experiments testing the robustness of our method to noise, and comparing primarily to ODR-BINDy [28].
@@ 4.5.1 Damped nonlinear oscillator
We considered increasing noise levels with fixed trajectory length and observation frequency for a nonlinear damped oscillator that was studied in [28]. The dynamics are given by
@@
We took the initial conditions $\boldsymbol{x}(0)=(2,0)$, and took samples of both variables with sampling period $\Delta t=0.05$ for $t\in[0,10]$. We first considered an example with noise level $\sigma=0.32$. We set trajectory regularization $\lambda=10^{-5}$, data weight $\alpha=10$, collocation penalty $\beta=10^{5}$, and used STLSQ with threshold $0.08$ and ridge parameter $100$. This is a surprisingly challenging example, especially with this level of noise–our model does not identify the correct support, but still performs quite well in estimating the state. We obtained $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=6.60\times 10^{-2}$, $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}=2.47\times 10^{-1}$, and the equations
@@
Next, we tested on a sequence of noise levels $\sigma=0.02,0.04,0.08,0.16,0.32,0.64$ and averaged the values of $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}$ and $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}$ over $32$ random seeds. In this case, we found that including ensembling [26] in the inner sparsifier greatly improved the consistency of the results for higher noise levels. We compared five models, weak SINDy[62], weak SINDy with ensembling, JSINDy, JSINDy with ensembling, and ODR-BINDy [26, 28]. We used PySINDy for ensembling and weak SINDy [45], and the GitHub repository [^2] https://github.com/llfung/ODR-BINDy associated to [28] for ODR-BINDy. We also tested with other two-step variants of SINDy, but they all performed worse than weak SINDy, so we omit them. As ODR-BINDy also produces state estimates, we compare the state estimation error between JSINDy, JSINDy with ensembling and ODR-BINDy.
@@
In terms of the coefficient error, and without ensembling the sparsifier, our model performs noticeably worse than ODR-BINDy; maximizing the Bayesian evidence seems to be a good model selection criterion. Upon introducing ensembling, our model selection stabilizes and the coefficient errors become comparable. In terms of state estimation, JSINDy both with and without ensembling performs well across varying levels of noise, regardless of the coefficient accuracy, owing to the stabilizing effect of the RKHS regularization and least squares relaxation.
@@
Figure 13: Results for the damped nonlinear oscillator across varying levels of noise. Left: coefficient error $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}$, Right: state estimation error $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}$. In this case, ensembling significantly improves the coefficient error for both weak SINDy and JSINDy. For JSINDy, including ensembling inside of the sparsifier brings the coefficient error from noticeably worse than ODR-BINDy to slightly better. Overall, JSINDy performs very well in state estimation across varying levels of noise.
@@ 4.5.2 Lorenz 63 example
We also tested the performance of our method on varying levels of noise and durations of observations with fixed observation frequency on the Lorenz 63 dynamics (40) with results in Figure 14. We applied JSINDy with data weight $\alpha=1$, collocation weight $\beta=10^{5}$, regularization $\lambda=10^{-5}$, and used SR3 as a sparsifier with an L1 penalty with weight 5, and relaxation coefficient $\nu=0.05$ [103]. We compare to ensemble weak SINDy and ODR-BINDy. The setup is identical to that of [28, Sec. 3.2], though we did not exclude the constant function from the library for weak SINDy. ODR-BINDy performed best in variable selection, likely owing to their approach of Bayesian evidence maximization. While also better than JSINDy in coefficient error, the difference is somewhat less pronounced. A much larger difference can be seen between JSINDy and ensemble weak SINDy, as well as the other models shown in [28, Sec. 3.2].
Figure 14: Experiment across different trajectory lengths and noise levels, averaged over 128 independent trials and $\Delta t=0.01$. Top: Proportion of trials where the correct support was exactly recovered Bottom: Relative coefficient error $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}$ of identified system.